# COADVISE: Covariate Adjustment with Variable Selection in Randomized Controlled Trials

### v3

@@ 4 Numerical Experiments
First, we generate 20 independent super-population full datasets, each containing $5\times 10^{7}$ independent units, with both potential outcomes $(Y(0),Y(1))$ available for each unit. We then compute the mean of $Y(1)-Y(0)$ across these 20 large datasets to obtain the true ATE. Given the large sample size and 20 replications, the associated uncertainty is negligible.
Second, we generate $M=500$ independent random sample datasets of size $N$, with $N$ varying between 100 and 500 to represent different sample size scenarios. In each Monte Carlo replication, the dataset consists of $N_{1}$ treated units and $N_{0}$ control units, with an expected 1:1 ratio such that $N_{0}+N_{1}=N$.
@@ 4.1 Data generating process
Continuous outcome: $Y(a)=50+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ where }e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a[3+\delta(\mathbf{X})])\}^{-1},$
for $a=0,1$, where $\epsilon\sim\mathcal{N}(0,1)$ i.i.d., and $\delta(\mathbf{X})$ characterizes the individual treatment effect. The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in our Online Supplemental Material for reproducibility.
@@ 4.2 Methods for comparison
All Include all $\mathbf{X}$ and $\mathbf{V}$, i.e., no covariate selection
Lasso Select covariates with non-zero regression coefficient estimates in Lasso regression (Bloniarz et al., 2016)
Adaptive Lasso Select covariates with non-zero regression coefficient estimates in adaptive Lasso regression (Zou, 2006)
Corr ($k$) Select the $k$ covariates with the highest marginal correlations with $Y$, where $k=1,3,10$ (Permutt, 1990)
Corr ($\xi$) Select covariates whose marginal correlations with $Y$ are greater than $\xi$, where $\xi=0.10,0.25$ (Permutt, 1990)
Pre-test ($\alpha$) Perform a univariate two-sample $t$-test on each covariate to test for differences between two groups at level $\alpha$. Select covariates where the test is rejected, with $\alpha=0.05,0.10$ (Senn, 1994; Zhao and Ding, 2023)
@@ 4.3 Performance criteria
  • Power%: The proportion of replications where the p-value is $\leq 0.05$ for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
@@ 4.4 Results
For brevity, we present the simulation results for the linear continuous outcome under both sample sizes of 100 and 500 in this section. Complete simulation results, including figures, are available in Appendix B of our Online Supplemental Material.
@@
Figure 2: Simulation results for the linear continuous outcome with sample sizes (a) $N=100$ and (b) $N=500$. In the Bias boxplots, the black dots represent the medians. In the CPplots, the red dashed line indicates the 95% coverage level, while the gray dashed lines indicate 93% and 97% coverage levels. In the Powerplots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power. Missing values in the CPand Powerfor ANHECOVA and ANCOVA are due to the non-obtainable sandwich variance estimators in high-dimensional settings.
@@
Figure 2 presents the results of biases via boxplots in the upper panels, and the results of CPand Powerin the lower four panels. We used different colors to distinguish between variable selection methods: green represents the estimator without variable selection (fully adjusted including all covariates), cyan represents Lasso and adaptive Lasso, gray indicates selection based on marginal correlation (as described in Table 2 for pre-specified $k$ or $\xi$), orange represents preliminary testing, and purple represents the simple estimator (which does not include any covariates).
@@
From Figure 2, we observe that when the sample size is small ($N=100$), there are many outliers in the boxplots for the ‘No selection” case for both AIPW and ANHECOVA estimators. This is expected, since the fully adjusted AIPW and ANHECOVA estimators include a relatively large number of covariates compared to the sample size. Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and $0.1$), and Corr ($k=1,3$) estimators exhibit wider bias box widths, with the Simple estimator widest, indicating that these methods are less efficient in this simulation setting.
@@
In terms of CPresults, we find that all methods are more stable around the 95% nominal level at the larger sample size, as the confidence intervals are Wald-type and rely on the large sample normal approximation. However, even when $N=500$, we observe below nominal CPfor both ANCOVA and ANHECOVA (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings). Additionally, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and $0.1$) estimators result in the lowest power. This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations. In contrast, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well across both sample sizes and all three estimators in terms of efficiency, variance estimation (as measured by CP%), and power.
Although we do not include the results of other DGPs, we briefly summarize them here. The results for the other continuous outcome are similar to those shown in Figure 2. For the two binary outcomes, we find that the efficiency gains observed in the boxplots are also similar to the continuous outcome cases, although the improvement relative to the Simple estimator is less pronounced due to the binary outcome being restricted to values in $\{0,1\}$. Moreover, the issue of non-obtainable sandwich variance estimators for ANHECOVA and ANCOVA is also present in the binary outcome cases.
@@ 5 Case Study
We apply the COADVISE framework to analyze the Best Apnea Interventions for Research (BestAIR) trial data, available through the National Sleep Research Resources (NSRR) (https://sleepdata.org/datasets/bestair/files/datasets). The dataset includes 169 participants, with 83 assigned to the active treatment group (continuous positive airway pressure [CPAP]) and 86 to a combined control arm. The primary objective of the study is to assess the impact of CPAP on 24-hour systolic blood pressure (SBP). The data includes 114 covariates, making it high-dimensional given the sample size ($N=169$).
Both the outcome and covariates contain missing values. Since only 25 observations are complete, we use the random forest method (Stekhoven and Bühlmann, 2012) (missForest option in our package) to impute all missing values. We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high ($\geq 130$ mmHg) or low ($<130$ mmHg) based on a 130 mmHg threshold.
@@ 6.2 Discussion
We acknowledge several limitations of the COADVISE framework. One notable concern is the issue of post-selection inference resulting from variable selection prior to ATE estimation (Fithian et al., 2014). Although our simulations suggest that our approach yields valid inferences, further theoretical investigation is warranted. In support of our method, we note that our model-based variance estimators (such as the sandwich or Delta methods) are robust to model misspecification, ensuring valid uncertainty quantification regardless of the variables selected for the outcome model(s). Furthermore, fitting the post-selection models using MLE or OLS improves efficiency, as demonstrated by the theory established in Section 3.2.
@@
Additionally, while our Coadvise R package includes a few commonly used missing data imputation methods for user convenience, we have not conducted a comprehensive investigation into their optimal integration within our framework, particularly in conjunction with variable selection. Future work could explore more robust approaches to handling missing data, including tailored imputation strategies that align with the underlying variable selection mechanisms to enhance both interpretability and statistical efficiency.
@@
Data and software availability
The BestAIR trial data used in Section 5 is available upon request at https://sleepdata.org. The R package Coadvise implementing the methodology of this paper is available at https://github.com/yiliu1998/Coadvise.
@@ Acknowledgments
The authors are grateful to Marlena Bannick for insightful discussions during the early stages of this research endeavor. Liu was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH) under Award Number T32HL079896. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
@@
S.1 R code for BestAIR trial study
@@ X The covariate vector/matrix/data frame $\checkmark$
Let $\mathcal{R}^{p}$ denote the $p$-dimensional Euclidean space $(p>0)$. Define $N_{a}=\sum_{i=1}^{N}I(A_{i}=a)$ as the number of units with treatment assignment $A_{i}=a$, $s_{N}^{2}(\mathcal{W})$ as the sample variance of the i.i.d. data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$, and $\text{cov}_{N}(\mathcal{W},\mathcal{V})$ as the sample covariance of the data $\mathcal{W}=\{W_{1},\dots,W_{N}\}$ and $\mathcal{V}=\{V_{1},\dots,V_{N}\}$. Similarly, $s_{N_{a}}^{2}(\mathcal{W})$ represents the sample variance for all $W_{i}\in\mathcal{W}$ with $A_{i}=a$, and $\text{cov}_{N_{a}}(\mathcal{W},\mathcal{V})$ denotes the sample covariance for all $W_{i}\in\mathcal{W}$ and $V_{i}\in\mathcal{V}$ with $A_{i}=a$, where $a=0,1$. Denote $\mathbb{P}_{N}[f(W)]=N^{-1}\sum_{i=1}^{N}f(W_{i})$ the empirical mean of any function of the i.i.d. observed data $\mathcal{W}$.
@@
S.3 Comparison of Coadvise and RobinCar
@@
rc.ancova <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
covariate_cols=covar.names, contrast_h = "diff", adj_method="ANCOVA")
@@
rc.anhecova <- robincar_linear(df=df.RobinCar, treat_col="A", response_col="Y",
contrast_h = "diff", covariate_cols=covar.names, adj_method="ANHECOVA")
@@
In this section, we present theoretical investigations into the efficiency gain achieved by the AIPW estimator for the ATE in RCT data. We adhere to the notation, setups, and assumptions established in the main text of the paper.
@@
Furthermore, this result can easily be extended situations where a variable selection step is performed before AIPW estimation. Suppose we select $s$ variables from $\mathbf{X}$ for $Y(a)$ model, where $0<s<p$. When fitting the OLS estimator, the only difference is that $\boldsymbol{\beta}\in\mathcal{R}^{s}$ instead of $\mathcal{R}^{p}$ for estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$. Here, $\widehat{\tau}_{\text{simple}}$ remains a special case of estimators in $\mathcal{G}^{\boldsymbol{\beta}}_{g}$. Therefore, the proof remains straightforward.
@@
Therefore, for any general $\mathbf{X}\not=\boldsymbol{0}$, we have $g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\not=\boldsymbol{0}$, since the logistic function satisfies $10<g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})<$ 1. Additionally, $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=\boldsymbol{0}$ only when $\boldsymbol{\beta}_{a}^{*}=\boldsymbol{0}$ (so $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=1/2$); for other $\boldsymbol{\beta}_{a}^{*}$, the sign of $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ can be either positive or negative. We now introduce Condition S.3.2 for using logistic regression to achieve efficiency gain.
@@
Figure S.3.1 illustrates the right-hand side of inequality in (ii). From the figure, if $\boldsymbol{\beta}_{a}^{*}$ is such that $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 1/2$, the (absolute value of) bias is allowed to be relatively large. However, when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ is either smaller or larger, the bias must be smaller, approaching zero when $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})\approx 0$ or $1$. In other words, when the true proportion of $Y(a)$ is very small or vary large, the model must better approximate the true value. But when the true proportion is around $50\%$, more error in the posited model is acceptable. The intuition is when the model $g(\mathbf{X}^{\prime}\boldsymbol{\beta}^{*})$ approaches a medium value, the uncertainty regarding whether the true $Y(a)$ is 00 or $1$ increases.
Figure S.3.1: An illustration for the bound of absolute bias in Condition S.3.2(ii) (right-hand side of the inequality). Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from $[0.11,\infty)$. When $g(\mathbf{X}^{\prime}\boldsymbol{\beta})=0.5$, the absolute bias is actually $\infty$; Panel (b) provides a further look for $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.3]$.
@@
Appendix S Additional Simulation Results
@@
where $\boldsymbol{I}_{50}$ is the $50\times 50$ identity matrix.
@@
Continuous outcome: $Y(a)=50+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ with }e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a[3+\delta(\mathbf{X})])\}^{-1},$

### v4

@@ 4 Numerical Experiments
[…]
Second, we generate $M=500$ independent random sample datasets of size $N$, with $N$ varying in {40, 100, 200, 500} to represent different sample size scenarios. The choice of 200 is is intended to mimic the sample size in the case study of Section 5, where $N=169$. In each Monte Carlo replication, the dataset consists of $N_{1}$ treated units and $N_{0}$ control units, with an expected 1:1 ratio such that $N_{0}+N_{1}=N$.
@@ 4.1 Data generating process
Continuous outcome: $Y(a)=30+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ where }$
$e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a\delta(\mathbf{X}))\}^{-1},$
for $a=0,1$, where $\epsilon\sim\mathcal{N}(0,1)$ i.i.d., and $\delta(\mathbf{X})$ characterizes the individual treatment effect. The details of $\boldsymbol{\beta}_{0}$, different cases of $\delta(\mathbf{X})$ (linear and nonlinear), and the distributions used for $\mathbf{X}$ and $\mathbf{V}$ can be found in our Online Supplemental Material (Section B) for reproducibility.
@@ 4.2 Methods for comparison
All
Include all $\mathbf{X}$ and $\mathbf{V}$, i.e., no covariate selection
Lasso
Select covariates with non-zero regression
coefficient estimates in Lasso regression (Bloniarz et al., 2016)
Adaptive Lasso
Select covariates with non-zero regression
coefficient estimates in adaptive Lasso regression (Zou, 2006)
Corr ($k$)
Select the $k$ covariates with the highest
marginal correlations with $Y$, where $k=1,3,10$ (Permutt, 1990)
Corr ($\xi$)
Select covariates whose marginal correlations
with $Y$ are greater than $\xi$, where $\xi=0.10,0.25$ (Permutt, 1990)
Pre-test ($\alpha$)
Perform a univariate two-sample $t$-test on each covariate to test
for differences between two groups at level $\alpha$. Select covariates
where the test is rejected, with $\alpha=0.05,0.10$
(Senn, 1994; Zhao and Ding, 2023)
@@ 4.3 Performance criteria
  • Power%: The proportion of replications where the p-value is 0.05$\leq$ 0.05≤ 0.05 for testing $H_{0}:\tau=0$ versus $H_{a}:\tau\neq 0$.
@@ 4.4 Results
For brevity, we present the simulation results for the nonlinear continuous and nonlinear outcomes under sample sizes of 40 and 200 in this section. Complete simulation results, including figures, are available in our Online Supplemental Material (Section B).
Nonlinear continuous outcome Refer to caption
Nonlinear binary outcome Refer to caption
@@
Figure 3: Simulation results for the nonlinear and binary continuous outcomes with sample sizes (a) $N=40$ and (b) $N=200$. In the Bias boxplots, the black dots represent the medians. In the CPplots, the red dashed line indicates the 95% coverage level, while the gray dashed lines indicate 93% and 97% coverage levels. In the Powerplots, the red dashed line indicates 100% power, and the gray dashed line indicates 80% power. Missing values in the CPand Powerfor ANHECOVA and ANCOVA are due to the non-obtainable sandwich variance estimators in high-dimensional settings.
Figure 4: Distributions of the AIPW estimates across 500 Monte Carlo replicates under nonlinear continuous and binary outcomes, with sample sizes $N=40$ and $N=200$. The red dashed lines represent the theoretical normal distributions based on the ATE truth and standard deviations of the 500 point estimates. The blue lines indicate the true ATE.
@@
Figure 3 presents the results of biases via boxplots in the upper panels, and the results of CPand Powerin the lower four panels under each outcome. We used different colors to distinguish between variable selection methods: green represents the estimator without variable selection (fully adjusted including all covariates), cyan represents Lasso and adaptive Lasso, gray indicates selection based on marginal correlation (as described in Table 2 for pre-specified $k$ or $\xi$), orange represents preliminary testing, and purple represents the simple estimator (which does not include any covariates).
@@
From Figure 3, we observe that when the sample size is very small ($N=40$), there are many outliers in the boxplots for the “No selection” case for all covariate-adjusted estimators. This is expected, since they include a relatively large number of covariates compared to the sample size. Across both sample sizes, the Simple, Pre-test (both $\alpha=0.05$ and 0.10.10.1) estimators exhibit wider bias box widths, with the Simple estimator widest. The Corr method does not exhibit consistent performance across different parameter choices of $k$ and $\xi$. For example, Corr with $k=1$ or $k=3$ is notably less efficient when the sample size is $N=200$, yet performs relatively well under $N=40$. In contrast, Corr with $\xi=0.10$ shows poor performance under $N=40$ but improves substantially with a larger sample size of $N=200$. Lasso and adaptive Lasso exhibit similar performance patterns under the same estimator and are overall the most efficient and consistently performing methods across simulation settings. An exception is observed with the ANHECOVA estimator under $N=40$ for both outcomes, where some outliers appear. Nevertheless, in all other scenarios—particularly when paired with the AIPW estimator—these methods consistently demonstrate stable gains in efficiency and statistical power. These findings suggest that Simple, Pre-test, and Corr methods are generally less efficient in our simulation setting, particularly when compared to Lasso and adaptive Lasso.
@@
In terms of CPand power, all methods exhibit greater stability around the 95% nominal level at the larger sample size ($N=200$), reflecting improved consistency of the variance estimators with increasing sample size. However, even when $N=200$, we observe below nominal or missing CPfor all three covariate-adjusted estimators (due to the inability to obtain HW sandwich variance estimators in high-dimensional settings). Interestingly, across both sample sizes, the Simple and Pre-test (both $\alpha=0.05$ and 0.10.10.1) estimators result in the lowest power. This finding for the Pre-test estimator is consistent with Zhao and Ding (2023), where they advise against its use in practice based on theory and simulations. Lasso and adaptive Lasso, when combined with all estimators, yield satisfactory CPunder the continuous outcome with $N=200$. However, under the binary outcome setting with $N=200$, they exhibit under-coverage when used with the AIPW estimator, though this performance improves when the sample size increases to $N=500$, as shown in our Online Supplemental Material. This empirical finding suggests that, for binary outcomes using logistic regression as the outcome model in AIPW, it is valuable to investigate more robust variance estimation methods for smaller sample sizes in future research. Overall, Lasso and adaptive Lasso selection methods combined with AIPW generally perform well (in terms of both point and variance estimations) across both sample sizes and both outcomes.
Futhermore, although we do not include the results of other DGPs, we briefly summarize them here. The results for other two outcomes (linear continuous and linear binary) are similar to those shown in Figure 3 and our comments above. At the largest sample size ($N=500$), the variance estimators following Lasso and adaptive Lasso selections perform more reliably with similar efficiency gains, yielding CPvalues that are closer to the 95% nominal level across all outcomes. However, the issue of non-obtainable sandwich variance estimators for ANHECOVA and ANCOVA without variable selection persists even at $N=500$. This indicates that, despite a seemingly adequate sample size relative to the data dimension, the sandwich variance estimator may remain computationally unstable or fail to converge when the number of covariates is large. This finding further highlights the advantage of applying variable selection to control the effective dimensionality, thereby facilitating the computation of model-based variance estimators.
Finally, to empirically assess the asymptotic normality of post-selection estimators following Lasso and adaptive Lasso, we present histograms of AIPW estimates after variable selection in Figure 4. Additional results for ANCOVA, ANHECOVA, and other simulation scenarios are provided in the Online Supplemental Material. As shown in Figure 4, the empirical distributions closely align with the corresponding theoretical normal densities (indicated by red dashed lines). As sample size increases, the estimates become more concentrated around the mean, and the agreement with the theoretical curves improves under both continuous and binary outcomes. Supplemental figures further support these findings, demonstrating strong normality even at moderate sample sizes (e.g., $N=100$ and $N=500$). Taken together, these results provide empirical evidence for the asymptotic normality of covariate-adjusted AIPW estimators following variable selection via Lasso or adaptive Lasso.
@@ 5 Case Study
We apply the COADVISE framework to analyze the Best Apnea Interventions for Research (BestAIR) trial data, available through the National Sleep Research Resources (NSRR) (https://sleepdata.org/datasets/bestair/files/datasets). The dataset includes 169 participants, with 83 assigned to the active treatment group (continuous positive airway pressure [CPAP]) and 86 to a combined control arm. The primary objective of the study is to assess the impact of CPAP on 24-hour systolic blood pressure (SBP). The data includes 114 covariates, making it high-dimensional given the sample size ($N=169$). We consider two outcomes: (i) continuous SBP values (in mmHg); and (ii) binary SBP values, classifying SBP as high (130$\geq$ 130≥ 130 mmHg) or low (<130<130< 130 mmHg) based on a 130 mmHg threshold.
Both the outcome and covariates contain missing values. Since only 25 observations are complete, we use the random forest method (Stekhoven and Bühlmann, 2012) (missForest option in our package) to impute all missing values. To justify the missingness pattern, we note that missForest assumes an MAR mechanism, which, while untestable in practice, is less restrictive than the MCAR assumption. For missing outcomes, we include all baseline covariates (e.g., age, gender, race, study site) in the imputation model, making the MAR assumption more plausible. Since treatment is randomized and all covariates are collected prior to treatment assignment, the outcome missingness mechanism is unlikely to depend on unobserved outcomes. Furthermore, missingness in baseline covariates is less concerning in randomized trials, as these covariates are independent of both treatment and outcome by design. In this application, which primarily aims to demonstrate the COADVISE framework, we impute all missing covariates under the MAR assumption and include them in the variable selection step to take full advantage of the high-dimensional structure of the data.
@@ 6.2 Discussion
We acknowledge several limitations of the COADVISE framework. First, we need to further investigate the post-selection inference arising from variable selection methods beyond Lasso and adaptive Lasso. Although our simulations suggest that all variable selection methods included in our package yield valid inferences under larger sample sizes, additional theoretical work is warranted to support these findings. In support of our method, we note that our model-based variance estimators (such as the sandwich or Delta methods) are robust to model misspecification, ensuring valid uncertainty quantification regardless of the variables selected for the outcome model(s). Furthermore, fitting the post-selection models using MLE or OLS improves efficiency, as demonstrated by the theory established in Section 3.2.
@@
Additionally, while our Coadvise R package includes several commonly used missing data imputation methods for user convenience, we have not undertaken a comprehensive investigation into their theoretical robustness or optimal integration within our framework, particularly in combination with variable selection. Future work could explore more robust imputation strategies that align with different variable selection mechanisms. We also plan to incorporate more advanced and recently developed imputation methods with theoretical guarantees that are compatible with our framework in future versions of the package.
@@
Data and code availability statement
The BestAIR trial data used in Section 5 is available upon reasonable request at https://sleepdata.org. The R package Coadvise implementing the methodology of this paper is available at https://github.com/yiliu1998/Coadvise.
@@ Acknowledgments
The authors are grateful to Marlena Bannick for insightful discussions during the early stages of this research endeavor. Yi Liu was supported by the National Heart, Lung, and Blood Institute (NHLBI) of the National Institutes of Health (NIH) under Award Number T32HL079896. This project is supported by the Food and Drug Administration (FDA) of the U.S. Department of Health and Human Services (HHS) as part of a financial assistance award U01FD007934 totaling \$1,674,013 over two years funded by FDA/HHS. It is also supported by the National Science Foundation under Award Number SES 2242776, totaling \$225,000 over three years. The contents are those of the authors and do not necessarily represent the official views of, nor an endorsement by, FDA/HHS, the NIH, or the U.S. Government.
@@
S.1 Summary of data imputation methods in COADVISE
Imputation method
Handles missing
covariates
Handles missing
outcome
Assumption
Complete-case analysis (cc) $\checkmark$ $\checkmark$ MCAR
Multiple imputation by chained equations
(mice; Van Buuren and Groothuis-Oudshoorn (2011))
$\checkmark$ $\checkmark$ MAR
Random forest
(missForest; Stekhoven and Bühlmann (2012))
$\checkmark$ $\checkmark$ MAR
Inverse probability weighting
(ipw; Robins et al. (1994))
$\times$ $\checkmark$ MAR
Missingness indicator
(missInd; Zhao et al. (2024);
Zhao and Ding (2022); Song et al. (2021))
$\checkmark$ $\times$ MAR
Table S.1: Summary of missing data imputation methods available in the current Coadvise R package.
S.2 R code for BestAIR trial study
@@ X The covariate vector/matrix/data frame $\checkmark$
[…]
@@
S.4 Comparison of Coadvise and RobinCar
@@
rc.ancova <- robincar_linear(df=df.RobinCar, treat_col="A",
response_col="Y", covariate_cols=covar.names,
contrast_h="diff", adj_method="ANCOVA")
@@
rc.anhecova <- robincar_linear(df=df.RobinCar, treat_col="A",
response_col="Y", contrast_h="diff",
covariate_cols=covar.names,
adj_method="ANHECOVA")
@@
S.5 A toy example for COADVISE with covariate-adaptive randomization
In this section, we present a toy example to illustrate the use of the COADVISE framework for covariate-adaptive randomization via the CoadviseCAR() function. As detailed in Section 3.4, our Coadvise package includes three estimators for ATE estimation: the simple unadjusted estimator (Simple), the stratified unadjusted estimator (Strata), and the AIPW estimator. For AIPW, all variable selection methods supported in the Coadvise() function are available.
Below, we provide the R code used to generate the toy example. We simulate data with $N=200$ observations and 100 covariates. Stratified randomization is performed across three strata—Low, Medium, and High—defined by a categorical variable. The potential outcomes are generated from a linear model involving the first three covariates, and the true ATE is set to $\tau=1$.
#### Data generation
n <- 200 # sample size
p <- 100 # number of covariates
strata_labels <- c("Low", "Medium", "High") # levels of strata
X <- matrix(rnorm(n*p), nrow=n, ncol=p) # covariates
colnames(X) <- paste0("X", 1:p)
A <- numeric(n)
# Generate treatment assignment with equal probability within each stratum
for (z in strata_labels) {
idx <- which(strata==z)
A[idx] <- rbinom(length(idx), 1, 0.5)
tau <- 1 # treatment effect (true ATE)
base <- X[,1] - 0.5*X[,2] + 0.3*X[,3]
Y0 <- base + rnorm(n) # potential outcome Y(0)
Y1 <- base + tau + rnorm(n) # potential outcome Y(1)
Y <- ifelse(A==1, Y1, Y0) # observed outcome
#### Analysis using CoadviseCAR() function and results
# Without variable selection for AIPW estimator
CoadviseCAR(Y=Y, A=A,
trt.name=1, ctrl.name=0, strata=strata,
X=X, conti.out=TRUE, var.sel.method="No")\$df.fit
method tau se ci.lwr ci.upr p
1 Simple 1.078399 0.2258623 0.6357165 1.521081 1.800795e-06
2 Strata 1.063202 0.2298618 0.6126817 1.513723 3.738767e-06
3 AIPW 3.796362 NaN NaN NaN NaN
# Lasso variable selection for AIPW estimator
CoadviseCAR(Y=Y, A=A,
trt.name=1, ctrl.name=0, strata=strata,
X=X, conti.out=TRUE, var.sel.method="Lasso")\$df.fit
method tau se ci.lwr ci.upr p
1 Simple 1.078399 0.2258623 0.6357165 1.521081 1.800795e-06
2 Strata 1.063202 0.2298618 0.6126817 1.513723 3.738767e-06
3 AIPW 1.085008 0.1412289 0.8082046 1.361812 1.554312e-14
From the analysis results, we observe that the Simple and Strata estimators yield similar point estimates and variance estimates. In contrast, the AIPW estimator without variable selection is highly biased and produces an unstable or unavailable variance estimate. However, after applying Lasso variable selection, the AIPW estimator provides an accurate point estimate with a smaller variance than both the Simple and Strata estimators, indicating a clear efficiency gain.
Online Supplemental Material
@@
In this section, we present our theoretical investigations into the efficiency gain achieved by the AIPW estimator for the ATE in RCT data. We adhere to the notation, setups, and assumptions established in the main text of the paper.
@@
[…]
@@
Therefore, for any general $\mathbf{X}\not=\boldsymbol{0}$, we have $g_{\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})\not=\boldsymbol{0}$, since the logistic function satisfies $0<g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})<1$. Additionally, $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=\boldsymbol{0}$ only when $\boldsymbol{\beta}_{a}^{*}=\boldsymbol{0}$ (so $g(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})=1/2$); for other $\boldsymbol{\beta}_{a}^{*}$, the sign of $g_{\boldsymbol{\beta}_{a}\boldsymbol{\beta}_{a}}(\mathbf{X}^{\prime}\boldsymbol{\beta}_{a}^{*})$ can be either positive or negative. We now introduce Condition S.3.2 for using logistic regression to achieve efficiency gain.
@@
[…]
Figure S.3.1: An illustration for the bound of absolute bias in Condition S.3.2(ii) (right-hand side of the inequality). Panel (a) provides the absolute biases over $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.9]$, which ranges from [0.11,)[0.11$\infty$)[ 0.11 , ∞ ). When $g(\mathbf{X}^{\prime}\boldsymbol{\beta})=0.5$, the absolute bias is actually $\infty$; Panel (b) provides a further look for $g(\mathbf{X}^{\prime}\boldsymbol{\beta})\in[0.1,0.3]$.
@@
Appendix S Additional Simulation Details
@@
where $\boldsymbol{I}_{50}$ is the $\times 50$ 50 identity matrix.
@@
Continuous outcome: $Y(a)=30+20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}+a\delta(\mathbf{X})+\epsilon,$
Binary outcome: $Y(a)\sim\text{Bern}(e(\mathbf{X},a)),\text{ with }e(\mathbf{X},a)=\{1+\exp(-20\mathbf{X}^{\prime}\boldsymbol{\beta}_{0}-a\delta(\mathbf{X}))\}^{-1},$